a 1300 N crate slides 14 m down the ramp that makes an angle of 30 degree with the horizontal. If the crate slides at a constant speed, how much thermal energy is created?
Answer #1

As the velocity remains constant.


Thermal energy produced = Potential energy lost
= 1300*(14sin30)
= 9100 J
Answer #2

As the velocity remains constant.
Thermal energy produced = Potential energy lost

= 1300*(15sin30)= 9750 J